Two moles of Helium gas (γ = 5/3) are initially at temperature 27oC and occupy a volume of 20 litres. The gas is first expanded at constant pressure until the volume is doubled. Then it undergoes an adiabatic change until the temperature returns to its initial value. What are the final volume?


(A)    113.13 × 10–3 m3                                     


(B)    213.13 × 10–3 m3  


(C)    313.13 × 10–3 m3                                   


(D)    13.13 × 10–3 m3 

Grade:12

2 Answers

vikas askiitian expert
13 years ago

pv = nRT

for first process pressure is kept constant so

v directionally proportional to temp

 Vi/Vf  =  Ti/Tf

 Vf = 2Vi ,  Vi = 20L  & Ti=300k

 so T2 = 600k

now second process is adiabatic so TVr-1 = c

                    Ti/Tf  =  (Vf/Vi)r-1

Ti = 600 ,Tf=300 , Vi=20*2=40L  & r=5/3 so

Vf = 80*21/2 L

    =113.13*10-3 m3

pranay -askiitians expert
13 years ago

thank you